OpenStudy (anonymous):

solve for x: ln(x^2 +1) - 3lnx= ln2?

OpenStudy (anonymous):

3lnx=ln(x^3) then by the propertie lna-lb=ln(a/b) ln((x^2+1)/x^3)=ln2 That means (x^2+1)/x^3 = 2 then solve that, x=1